SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Aug 2, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Yaw Ming Chia, Al Eisner, Fang Fang, Selina Li, Brian Lindquist, Dong Su, Dan Walker.

Operations

  • PEP tuned aggressively and pushed for getting high luminosity but background was very high as well. We got paged on Friday evening because of the high deadtime (~10% or so deadtime), but DCH had the biggest deadtime. Selina showed some plots for looking at the L1 and L3 rate, lumi, hadron A, B, and CMS etc. in a strip chart for Run 5, and for the month of July. A plot of L1 rate vs lumi for a comparison of Run 4 and Run5 was also shown.

EMT

  • Dan is performing test of the prescale for EMT with his own database. This should be completed in a few days. He might do a test with the EMT test stand, but will need help from Tom to get the test stand setup to do the test. Su Dong suggested to use the EMT fastmon to monitor the bit.
  • Dan showed some turn-on curves for the EMT fastmon. We are happy with the EMT M turn-on for mu-mu, matched L3 track.

DCT/DCZ

  • Ming talked about the problem with the TSFi (prototype) board at test stand. See the plots for the DCH-DCT calibration shown. The troublesome TSFi board was replaced with a spare, and now the test stand finallly has a working set of hardware and Ming can start his real testing for the investigation of power-cycling problem with the TSF crate.
  • There was discrepancy for the TSF->BLT data and simulation. Need to check if the simulation is doing what data are doing.
  • The new online release was installed at IR2 today. Fang and Selina linked in the new ROM lib (libL1DctCalApp.o) for production. We need to take a run to test this. Rainer asked if the TSF and BLT/ZPD are using the same ROM lib. Why do we still have to make a ROM lib built and link this in by hand whenever there is an online release going to installed at IR2? Is possible to have this automatically updated with the official release? We should aim to take this dependency out so that we don't have to link by hand even though there is no code changes for us.
  • It's decided that we will have the old DCT on for November. However, we still haven't reached a conclusion on whether we (1)turn it off now and turn it back on in November, or (2)keep running with it until Nov. There was discussion of things need to be done if it is the first case. A list of things we can think of are: to create a new configuration key with the old DCT crate not included; to disable some of the sequences related to the old DCT in fastmon; to comment out the lines for creating histograms for the old DCT system for livefastmon .ps archives. If we continue running with both the new and old DCT systems, then no extra work is needed to carry out.

GLT

  • Su Dong has the GLT FC PROM ready to test at IR2 around 4pm today. The plan is to take cosmics for a validation test today since PEP is having MD till midnight or so. And then take collision data with the new PROM tomorrow.

L3

  • Last week, Rainer implemented a fix for the L3TPCA server that should cure the occasional glitches causing alarms and "damage reports" at the start of a new run.

Data Quality and Monitoring

  • We decided that run 21477 is unusable due to the big gap in phi for L3 tracking (it was about 50% inefficient in that region). This run was about 3 pb^(-1). In this run, the OPR plots for the efficiency of the TSF looked okay, but the inefficiency showed up in L3 tracking. It was suggested to look at the TSF occupany plots instead, but these weren't available for run2, so we could run a fastmon job on this run to confirm.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.